Solve the system of equations using the substitution method.
3x + y = 2
2x – y = 3
This is what I cam up with, but It doesn't seem correct.
3x + y =2
Isolate the y by sub. 3x from both sides
3x-3x + y =3x +2
substitute 3x + 2 for y
2x - 3x + 2=3
This is where I get lost, not sure if this is even correct

Answer by Nate(3500)

You know what you're doing, but you have to look it over really well...
Lets look at your problem:
3x-3x + y =3x +2 ...... the 3x should be negative because you subtracted it
-------------------------------------------
Know, lets begin:
3x + y = 2 original equation
y= -3x + 2 sub. 3x
-------------------------------------------
2x – y = 3 original equation
2x - (-3x + 2) = 3 substitute
2x + 3x - 2 = 3 distribute the negative one
5x = 5 simplify
x=1
-------------------------------------------
know, lets plug 1 for x in one of the equations
3x + y = 2 orginal equation
3 + y = 2
y= -1